OpenStudy (anonymous):

Solve the following inequality and write your solution using set builder notation. 5/9(x) >= - 3/4 need help to solve please

OpenStudy (anonymous):

5/9(x) >= - 3/4 5/9-5/9(x) >= - 3/4 -5/9 x>= -3/4(9)-5/9(4) x>= -27/36 - 20/36 x>=-7/36